In the process of setting up some cabling at 1Y2, I must've bumped a cable to the c1lsc expansion chassis. Anyways, the c1lsc models crashed. I ran the reboot script around 530pm PDT. Usual locking behavior was recovered after this. The work at 1Y2 was:
The IN2 to CM board was already connected to I single ended output of the ALS X demodulator. The ~100 Hz UGF digital locking using the CM_SLOW path is straightforward but I didn't have any success with the AO path tonight. I wonder how high BW this lock can be made without injecting a ton of noise into the IMC loop, given that the EX uPDH only has ~ 10 kHz UGF.
Attachment #1 shows the spectra of the ALS signal
Attachment #2 is an OLTF measurement.